OpenStudy (anonymous):

T_T Please help me. Solve the equation. Check for extraneous solutions. [(x-3)/(x+2)] = [(x-1)/(3x-1)] Please show work! and thanks.

OpenStudy (anonymous):

I can do these all day.... Just look at what I said in my last thing. Review the book...

OpenStudy (anonymous):

Yeah i have and i just don't understand it! But feels like i'm so close to the answer!

OpenStudy (anonymous):

I could show my work on here but it would look so clear... : /

OpenStudy (anonymous):

x=1/2 x=5

OpenStudy (anonymous):

wouldnt*

OpenStudy (anonymous):

idk how to get there...

OpenStudy (phi):

Do they teach you to cross multiply?

OpenStudy (anonymous):

Yes, but i tried that... :L I'll try again.

OpenStudy (phi):

the problem should turn into (x-3)(3x-1) = (x-1)(x+2) do it one side at a time (x-3)(3x-1) = ??

OpenStudy (phi):

If you need more info, watch this http://www.khanacademy.org/math/algebra/polynomials/v/multiplication-of-polynomials

OpenStudy (precal):